Find C(n, x)p*q* for the given values of n, x, and p. (Round your answer to four decimal places.) n=8, x=4,p= 4 Need Help? Read It + -/1 points TanFin 12 8.4.012. Use the formula C(n, x)p*qX to determine the probability of the given event. (Round your answer to four decimal places.) The probability of exactly five successes in ten trials of a binomial experiment in which p = 2 Need Help? Read It Answer Both Questions, please. Thank you
Answer #1

A) n = 8, x = 4, p = 1/4 = 0.25

P(X = 4) = 8C4 * (0.25)^4 * (0.75)^4 = 0.0865

B) n = 10, x = 5, P = 1/2 = 0.5

P(X = 5) = 10C5 * (0.5)^5 * (0.5)^5 = 0.2461